will you show me how to work these?
Pietleas to do tof taofou H+NO HAO .2M .36M rem Salecfon of KNo CoMes K-4.Sx10 NHOH NH,0 WO add of NH C OH K 1.8x10 -S
Answer #1

A) it forms acidic buffer,

pH = pKa + log([NO2-/HNO2])

= 3.35+log(0.21/0.36)

= 3.11

B) it basic buffer,

pOH = pKb + log(NH4Cl/NH4OH)

= 4.74+ log(0.45/0.60)

= 4.62

[OH-] = 2.46×10-5 M

pH = 14-4.62

= 9.38